# In-context KV-Cache Eviction for LLMs via Attention-Gate

### v2

@@ A.1 Results of Continual Pre-training on Mistral
We conducted continual pre-training on Mistral using 5,000 samples from RedPajama (Computer, 2023), and the results are shown in Table 4. Compared to the performance of Llama2-7B presented in Table 1, Mistral’s performance slightly declined. We hypothesize that this may be due to the distribution of RedPajama’s data being less suited to Mistral. Additionally, this raises the question of whether KV-Cache eviction is model-dependent, and whether its effectiveness is related to the model’s expressive power. Although the parameter counts of Mistral-7B and Llama2-7B are similar, Mistral-7B significantly outperforms Llama2-7B. This could suggest that Mistral is utilizing more tokens or scoring them with finer granularity, which results in fewer redundant tokens and thus makes eviction less effective.
@@
A.2 More Visualization
Figure 4: The complete version of Figure 3.
Figure 5: More samples for the same scenario of Figure 3.
Figure 4 provides a comprehensive view of the layers and attention heads from Figure 3. Additionally, four sample scenarios for the same setup are presented in Figure 5.
A.3 More Ablation
@@
%Eviction 65.66 44.48 65.14 24.28 68.18 63.44 55.20
In this section, we present additional ablation results in Table 5.

### v3

@@ A.2 Results of Continual Pre-training on Mistral
We conducted continual pre-training on Mistral-7B (Jiang et al., 2023) using 5,000 samples from RedPajama (Computer, 2023), and the results are shown in Table 5. Compared to the performance of Llama2-7B presented in Table 4, Mistral’s performance slightly declined. We hypothesize that this may be due to the distribution of RedPajama’s data being less suited to Mistral. Additionally, this raises the question of whether KV-Cache eviction is model-dependent, and whether its effectiveness is related to the model’s expressive power. Although the parameter counts of Mistral-7B and Llama2-7B are similar, Mistral-7B significantly outperforms Llama2-7B. This could suggest that Mistral is utilizing more tokens or scoring them with finer granularity, which results in fewer redundant tokens and thus makes eviction less effective. Furthermore, it is possible that Mistral’s use of grouped-query attention (GQA), which inherently involves compression, may make it more challenging to increase the eviction ratio effectively in this context.
@@
A.3 More Visualization
Figure 5 provides a comprehensive view of the layers and attention heads from Figure 4.
Figure 5: The complete version of Figure 4. Two key observations emerge in (i): 1. the first two layers are denser compared to the subsequent layers, and 2. bright-yellow vertical lines, representing critical tokens for inference, consistently appear across heads in deeper layers.
A.4 More Ablation
In this section, we present additional ablation results in Table 6.
@@ A.4 More Ablation
%Evict. 65.66 44.48 65.14 24.28 68.18 63.44 55.20